Pentane (C5H12) burns according to the following unbalanced reaction:

____C5H12 (g) + ____O2 (g) → ____CO2 (g) + ____H2­O(g)

What volume of O2 is required to produce 38.0 L of Carbon dioxide? Assume STP.
1
Expert's answer
2017-04-21T12:33:05-0400
1. Write the balanced equation:
C5H12 (g) + 8O2 (g) → 5CO2 (g) + 6H2O(g)
The equation shows that 8 moles of oxygen are required to produce 5 moles of carbon dioxide.

2. Convert moles to liters. 1 mole of any gas at STP has a volume of 22.4 L.
So, from (1) we can say that (8 * 22.4) = 179.2 L of oxygen are required to produce (5 * 22.4) = 112.0 L of carbon dioxide.
Therefore in order to produce 38.0 L of carbon dioxide we need (38.0 * 179.2) / 112.0 = 60.8 L of oxygen.



Answer:
60.8 liters of oxygen are needed to produce 38.0 liters of carbon dioxide.